A rectangle is four times as long as it is wide. If it has an area of 9 square inches, what are its dimensions?
Answer by unlockmath(1688)

Hello,
Let's have x be the width and 4x be the length. Now an equation can be set up as:
x(4x)=9 Rewritten as:
4x^2=9 Divide wach side by 4 to get:
x^2=9/4 Square root each side to result in:
x=3/2 or 1 1/2 inches is the width and length is:
6 inches.
There we go. Multiply these two answers tosee if it comes up with 9 sq inches.
RJ
There is a book that might help which I wrote:
www.math-unlock.com